Find the complement rate for each discount in the given series of trade discounts. Then find the net price, using the complement method. Do not round intermediate computations. Complement Rates List Price Trade Discount First Second Third Net Price $4,000 60%, 30%, 20%

First %

Second %

Third %

Net Price $
Answer #1
List Price = $       4,000
Less: Trade Discount:
First Discount @60% = $       2,400
Second Discount @30% = $           480
Third Discount @20% = $           224
Net Price = $           896
Workings:
List Price = $       4,000
Less: First Discount @60% ($4000*0.6) = $       2,400
Net Price = $       1,600
Less: Second Discount @30% ($1600*0.3) = $           480
Net Price = $       1,120
Less: Third Discount @20% ($1120*0.2) = $           224
Net Price = $           896